OpenStudy (anonymous):

The strength of a magnetic field follows _____. an inverse square law the law of domains Coulomb's law the law of poles

OpenStudy (anonymous):

Will give medals!(:

OpenStudy (anonymous):

Hint: The magnetic field lines for infinitely long straight wire \[B = \frac{ \mu _0 I }{ 2 \pi } \times \frac{ 1 }{ r }\]